OpenStudy (anonymous):

what is the solution of the system y=-3x, 3x+2y=6. if someone answers i will give a medal.

OpenStudy (anonymous):

how long is the answer?

OpenStudy (anonymous):

Plug y=-3x in 3x+2y=6. 3x+2(-3x)=6 3x+(-6x)=6 -3x=6 x=-2 Now, plug x=-2 in either equation to find the value of y. Good luck

OpenStudy (anonymous):

oh that makes sense thanks.